The debt-to-equity ratio of the CI Corp. is 0.5. To calculate the CI Corp’s cost of equity, use the following information: The CI Corp’s beta is 1.2, expected return on the market is 8%, and the risk-free rate is 2%. Its pretax cost of debt is 4%, what is the company’s WACC? The tax rate is 35%
Answer #1

WACC = Weight of debt * Cost of debt * (1 - tax) + Weight of equity * Cost of Equity

Debt-Equity ratio = 0.5

=> If Equity = 1, Debt = 0.5

Weight of Debt = Debt/(Debt + Equity) = 0.5/(1+0.5) = 1/3

Weight of Equity = 1 - Weight of debt = 1- 1/3 = 2/3

Based on CAPM,

Cost of Equity = Risk free rate + Beta * (Expected market return - Risk free rate)

Cost of Equity = 2% + 1.2 * (8% - 2%) = 9.20%

WACC = (1/3) * 4% * (1 - 35%) + (2/3) * 9.20%

WACC = 7.00%